OpenStudy (anonymous):

A student opens a mathematics book to two facing pages. The product of the pages is 1806. The first page number is__? The second page number is __?

OpenStudy (anonymous):

x -y = 1 xy = 1806 x(x-1) = 1806 x^2 - x -1806 = 0 x = 42 first page = 42 secnd = 43

OpenStudy (anonymous):

42*43 = 1806

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

yw